A container has volume 5700 cubic meter. it takes 12 hrs to drain this container. it is filled with water. What is the rate of mass flow of water in kg/s when the density of water is 1g/cm3

Grade:12

1 Answers

ANSHU GOYAL
12 years ago

Density of water is 1000kg/m3.

therefore mass of given water in kg is = density * volume= 1000*5700.

it takes 12*60*60 sec to drain this mass of water.

therefore rate of mass flow = 5700000/43200=131.94 kg/s